OpenStudy (anonymous):

Which pair of ratios forms a proportion? A. 11/36, 121/180 B. 13/21, 39/63 C. 14/23, 70/92 D. 20/51, 4/5

OpenStudy (anonymous):

"forms a proportion" means that they are equal fractions, or equal ratios. Here's an example. 1/2, 2/4

OpenStudy (anonymous):

ohhh, so then the answer would be A right?

OpenStudy (anonymous):

I do have another question, would you mind helping me with it?

OpenStudy (anonymous):

Which ratio forms a proportion with 4/5 A. 2/3 B. 4/10 C. 16/20 D. 16/25 Could you show me how you would figure this out?

OpenStudy (anonymous):

There are two good ways to compare those fractions and see if they're equal. If you're allowed to, you can put them in your calculator to see what they look like as a decimal. Otherwise, rewrite them with a common denominator. Here's what that looks like for A: Method 1) 11/36, 121/180 11/36 = 0.305 121/180 = 0.672 Equal? No. Method 2) 36 and 180 have common denominator of 180 because 36 *5 = 180 11/36 *(5/5) = 55/180 121/180 = 55/180? No.

OpenStudy (anonymous):

Don't move on to your second question yet. Let's make sure you understand the first =)

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Read that explanation and ask me questions. Do you understand what I did to check choice A? Do you understand why?

OpenStudy (anonymous):

okay, I understand, the problem I tend to have though is with factoring, im not that good at it. So, when using the second method, I tend to take much more time to solve. Factoring is the way I learned how to solve these, I have never seen the other one, im gonna use it to solve the first question, ill let you know when I have my answer... :)

OpenStudy (anonymous):

Okay =)

OpenStudy (anonymous):

but how do I convert the fraction to a decimal on the calculator? im new with using an algebraic calculator...

OpenStudy (anonymous):

okay, I figured it out, and the answer is B :)

OpenStudy (anonymous):

are you still there?